OpenStudy (joshoyen):

Factor the expression. 2x^2 - 6x + 4

OpenStudy (md152727):

Have you tried to solve that?

OpenStudy (anonymous):

2(x^2-3x+2

OpenStudy (joshoyen):

Just confused on the 2x^2 cus I usually solve without a number in front of x.

OpenStudy (tkhunny):

You'll have to get used to that. For now, listen to @sSoyanii Factor it out of the entire expression.

OpenStudy (joshoyen):

So its 2(x-1x-2x+2)

OpenStudy (joshoyen):

How do I do the next step

OpenStudy (tkhunny):

What is that? 2x^2 - 6x + 4 = 2(x^2 - 3x + 2) This is the Distributive Property of Multiplication over Addition. Factor the rest: 2(x^2 - 3x + 2) = 2(x _____ _____)(x ______ _____)

OpenStudy (joshoyen):

2(x-1)(x-2)

OpenStudy (joshoyen):

How do I factor this one. 2x^2 -11x + 5

OpenStudy (tkhunny):

I think you have it. Factor everything the same way. Look for common factors and factor them out. Setup what you can set up. 2x^2 - 11x + 5 = (2x ____ ____)(x ____ ____) The two end blanks must be filled in with factors of 5. 1*5 that's it. The two middle blanks must be the same sign, since 5 is positive. Just keep this in mind: (ax+b)(cx+d) = acx + (ad+bc)x + bd In your problem 1) ac = 2 2) bd = +5 3) (ad + bc) = -11

OpenStudy (joshoyen):

thank you!